# New Methods for Network Count Time Series

### v1

@@ 1.1 Network time series
A network time series is the pair [$\left[\{X_{t}\}_{t},G\right]$, where $G=(V,E)$ is a graph (or network), where $V$ is a set of vertices or nodes, with $|V|=N$, and edge set $E$ and $\{X_{t}\}_{t}$ is a $N$-dimensional multivariate time series $X_{i,t}$ for $i\in V$ and times $t=1,\ldots,T$ for some integer $T>0$. We use the notation $i\leftrightsquigarrow j$, if $i\in V$ is directly connected to $j\in V$. Given a subset of nodes $A\subset V$, then the neighbourhood set of $A$ is defined by
@@ 1.2 GNAR models
A particular recent network time series model is the generalized network autoregressive model (GNAR) introduced by Knight et al. (2016), see also Zhu et al. (2017) and Knight et al. (2020). The GNAR model assumes that each node (variable) of the multivariate time series is influenced by an standard autoregressive term and contributions from neighbours in the network at earlier times. A GNAR$(p,[s_{1},\ldots,s_{p}])$ model is given by:
@@
The components are the GNAR model in (3) are 1. the autoregressive parameters, $\{\alpha_{i,j}\}_{i=1,\ldots,N;j=1,\ldots p}$ explain how past values of $X_{i,\cdot}$ contribute to $X_{i,t}$. Every vertex (or node) in the graph has their own autoregressive sequence. For some data sets, it is possible to set $\alpha_{i,j}=\alpha_{j}$. In other words, a common $\{\alpha_{j}\}_{j=1,\ldots,p}$ applies to each vertex and the model is then called a global GNAR process. 2. The term $\sum_{q\in{\mathcal{N}}^{(r)}(i)}\omega_{i,q,c}^{(t)}X_{q,t-j}$ elicits a contribution from each $r$th stage neighbour, $q$, of vertex $i$, lagged by time $j$ relating to covariate type $c$. The $w^{(t)}_{i,q,c}$ are called connection weights, which are often related to the local positioning of neighbours of vertex $i$ relative to $i$ and each other. The weights could be inverse distance weights, see Knight et al. (2020) for a definition and an example of how they are used. 3. The $s_{j}$ control the maximum number of stages of neighbours at lag $j$ for node $i$.
@@ 1.3 Features of GNAR models
Once formulated and model selection completed, GNAR models can be fit efficiently and rapidly by using existing linear modelling software. Knight et al. (2020) and Nason and Wei (2022) demonstrate that well-fitting GNAR-based models are highly parsimonious and also deal with missing data well.
@@ 1.4 Some univariate count time series models
We briefly review three types of popular time series models: the INAR, GAR and NAR models and some closely related ones. See Davis et al. (2021) for a comprehensive recent review.
The INteger-valued AutoRegressive model is based on thinning operations. The INAR$(1)$ model was introduced by Al-Osh and Alzaid (1987) and the general INAR$(p)$ model by Jin-Guan and Yuan (1991). The binomial thinning operation is defined as follows. Let $Y_{1},Y_{2},\ldots$ be a collection of independent and identically distributed Bernoulli random variables with (probability of success) parameter $q\in(0,1)$. Let $X$ be a non-negative integer random variable. The binomially-thinned random variable $Y$ is given by
@@
Let $\{\alpha_{j}\in(0,1)\}_{j=1}^{p}$, then the univariate INAR$(p)$ model for count time series, $X_{t}$ is given by
$X_{t}=\sum_{i=1}^{p}\alpha_{j}\circ X_{t-j}+\epsilon_{t}.$ (6)
Here $X_{t}\in{\mathbb{N}}_{0}$, the set of non-negative integers, and $\epsilon_{t}\in{\mathbb{N}}_{0}$ is a set of uncorrelated random variables. If the $\{\epsilon_{t}\}$ are Poisson-distributed, then $X_{t}$ is then called a Poisson-INAR$(p)$ process.
[…]
@@
The generalized autoregressive GAR$(p)$ model, without covariates, is characterised by the following ‘mean-relation’
$g(\mu_{t})=\sum_{i=1}^{p}\alpha_{j}\mathcal{A}(X_{t-j}),$ (7)
where $\mu_{t}={\mathbb{E}}(X_{t})$, $g$ is a link function, $\alpha_{j}\in{\mathbb{R}}$ and $\mathcal{A}$ is some function that modifies the autoregressive relations. GAR processes are related to generalised linear models. GAR$(p)$ models are special cases of the GARMA$(p,q)$ model as introduced by Benjamin et al. (2003). Often, if $X_{t}\sim\text{GARMA}(p,q)$ given past history follows some exponential family, and popular choices for the conditional distribution are Poisson, binomial and gamma. Many popular nonlinear models extended from the integer-valued generalized autoregressive conditional heteroskedasticity (INGARCH) model use the same link function idea as GARMA. For example, the log-linear Poisson autoregression from Fokianos and Tjøstheim (2011) or the $\operatorname{softplus}$-INGARCH model from Weiß et al. (2020).
The nonlinear autoregressive (NAR) process Jones (1978) has similarities to GAR — they both involve a link or response function as follows:
@@ 1.5 Poisson network autoregression
The first network time series model for count data was the Poisson network autoregression (PNAR) and Poisson GNAR, which are a count-valued network time series models introduced by Armillotta and Fokianos (2024) based on the network autoregression models from Zhu et al. (2017) and Knight et al. (2016, 2020), respectively. The linear PNAR$(p)$ model assumes for vertex $i$ and time $t$ that $X_{i,t}\sim\operatorname{Poi}(\lambda_{i,t})$, where
@@
The PNAR model permits interdependence among nodes at time $t$ and this correlation is induced via copula methods, which depends on unknown parameters in addition to the $\alpha$s and $\beta$s above. Interesting conditions for stationarity and ergodicity for PNAR$(p)$ are
@@
Armillotta and Fokianos (2024) further present a count data extension of the GNAR model Knight et al. (2016, 2020) termed the Poisson GNAR where the conditional mean is based on GNAR components as
@@
Armillotta and Fokianos (2024) also consider a log-linear PNAR model $X_{i,t}\sim\operatorname{Poi}\{\exp(\nu_{i,t})\}$ where the linear predictor, $\nu_{i,t}$, is identical to (9) except that the $X$ terms are replaced by $\log(X+1)$ and the $\alpha$ and $\beta$ parameters can be real numbers. The $+1$ in the $\log$ term in the linear predictor is to handle zero values of $X$. Parameter estimation for both the linear and log-linear PNAR models is performed by quasi-maximised likelihood estimation (QMLE).
@@ 2.1.1 GNARI model definition
where $\circ$ denotes thinning as before, $\alpha_{i,j},\beta_{j,r}\in[0,1]$, $\epsilon_{i,t}$ is assumed to be non-negative independently distributed for each node $i$ and time $t$, and identically distributed for the same $i$, i.e. for each $i,t$ we have ${\mathbb{E}}(\epsilon_{i,t})=\lambda_{i}$ for some $\lambda_{i}\in{\mathbb{R}}^{+}$. This specification permits us to assign a different mean and variance for each node $i$, but we can choose the make the process ‘global’, i.e. $\lambda_{i}=\lambda$ as with the ‘global-$\alpha$ specification of the original GNAR processes. As with INAR processes, we often assume the $\epsilon_{i,t}$ to be Poisson-distributed.
GNARI models share the same limitation of not permitting negative correlations as INAR. However, they are a popular model in the regular time series case and worth study. Parameter estimation can be carried out by conditional least-squares or conditional maximum likelihood, which we develop next.
@@ 2.1.2 GNARI conditional distribution and stationarity
We introduce the additional notation $Y^{(r)}_{i,j}$ to be the contribution of the $r$th-stage neighbours of node $i$ that are $j$ time steps prior to time $t$, i.e.
@@
where $B_{j,r,k}$ are Bernoulli$(\beta_{j,r})$ random variables. The $Z_{i,j,t}^{(r)}$ quantity encapsulates the full second term in (12).
@@ 2.1.3 GNARI process autocovariance
We now derive the autocovariance function $\Gamma(h)=\operatorname{Cov}(\mathbf{X}_{t},\mathbf{X}_{t-h})$ for (17), under stationarity. From the proof of Lemma 2.2, a GNARI($p,[s_{1},\dots,s_{p}]$) process is equivalent to the MGINAR(1) process as defined in (49). Then, by Latour (1997) Section 4, the autocovariance function for (49), $\Gamma^{\prime}(h)$, satisfies
@@
where $B$ is the variance matrix corresponding to the thinning operation $A\circ\cdot$, $\mu_{Y}=\mathbb{E}[\mathbf{Y}_{t}]$, and $\Sigma_{e}=\operatorname{Var}[\mathbf{e}_{t}]$ and $\mathbf{Y}_{t}$ is defined in (49). For GNARI we have $B=(1-A)A$ that is
$B=\begin{bmatrix}A_{1}(1-A_{1})&A_{2}(1-A_{2})&\dots&A_{p-1}(1-A_{p-1})&A_{p}(1-A_{p})\\ 0&0&\dots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\dots&0&0\end{bmatrix},$ (21)
$\mu_{Y}=(\mu_{X}^{T},\ldots,\mu_{X}^{T})$ and $\Sigma_{e}=\operatorname{diag}(\lambda_{1},\ldots,\lambda_{N},0,\ldots,0)$. It is easy to show that the autocovariance function for (17), $\Gamma(h+j)$, is the $[jN+1:(j+1)N,jN+1:(j+1)N]$ submatrix of $\Gamma^{\prime}(h)$. More specifically, the autocovariance function $\Gamma^{\prime}(h)$ can be written in terms of $\Gamma(h+j)$ as
@@ 2.1.4 Conditional least squares estimation
Let $\mathbf{\beta}=(\alpha_{1},\beta_{1,1},\ldots,\beta_{1,s_{1}},\ldots,\alpha_{p},\beta_{p,1},\ldots,\beta_{p,s_{p}},\lambda)^{T}$ be the parameter of interest, $\mathcal{F}_{t}$ be the $\sigma$-algebra generated by $\mathbf{X}_{t},\mathbf{X}_{t-1},\ldots$, then the conditional least squares estimator $\hat{\mathbf{\beta}}_{n}$ minimizes
$Q_{n}(\mathbf{\beta})$ $=\sum_{t=1}^{n}||X_{t}-\mathbb{E}(X_{t}|\mathcal{F}_{t-1})||^{2}$ (23)
@@
where $S_{t,i,r,j}=\sum_{q\in\mathcal{N}^{r}_{t}(i)}w_{i,q}X_{q,t-j}$. As with standard least squares, the estimator $\hat{\mathbf{\beta}}_{n}$ can be obtained by solving the usual normal equations: $X^{T}X\mathbf{\beta}=X^{T}\mathbf{Y}$.
@@ 2.1.5 Asymptotic properties
Let $\hat{\mathbf{X}}_{t|t-1}(\mathbf{\beta})=\mathbb{E}_{\mathbf{\beta}}(\mathbf{X}_{t}|\mathcal{F}_{t-1})$,
$f_{t|t-1}(\mathbf{\beta})=\mathbb{E}[\{\mathbf{X}_{t}-\hat{\mathbf{X}}_{t|t-1}(\mathbf{\beta})\}\{\mathbf{X}_{t}-\hat{\mathbf{X}}_{t|t-1}(\mathbf{\beta})\}^{T}|\mathcal{F}_{t-1}]$ (28)
and let $\mathbf{\beta}_{0}$ be the true value of $\mathbf{\beta}$. Here $\mathbf{X}_{t}=(X_{1,t},X_{2,t},\ldots,X_{p,t})$ is to be considered as a row vector.
Assuming GNARI process stationarity, Theorem 3.2 of Tjøstheim (1986) implies that, under the regularity conditions proposed in that theorem
$\sqrt{n}(\hat{\mathbf{\beta}}_{n}-\mathbf{\beta}_{0})\xrightarrow{}\operatorname{MVN}(0,U^{-1}RU^{-1}),$ (29)
@@
$U=\mathbb{E}\left\{\frac{\partial\hat{\mathbf{X}}_{t|t-1}^{T}}{\partial\mathbf{\beta}}(\mathbf{\beta}_{0})\frac{\partial\hat{\mathbf{X}}_{t|t-1}}{\partial\mathbf{\beta}}(\mathbf{\beta}_{0})\right\},$ (30)
@@
$R=\mathbb{E}\left\{\frac{\partial\hat{\mathbf{X}}_{t|t-1}^{T}}{\partial\mathbf{\beta}}(\mathbf{\beta}_{0})f_{t|t-1}(\mathbf{\beta}_{0})\frac{\partial\hat{\mathbf{X}}_{t|t-1}}{\partial\mathbf{\beta}}(\mathbf{\beta}_{0})\right\}.$ (31)

### v2

@@ 1.1 Network time series
A network time series is the pair $\left[\{X_{t}\}_{t},G\right]$, where $G=(V,E)$ is a graph (or network), where $V$ is a set of vertices or nodes, with $|V|=N$, and edge set $E$ and $\{X_{t}\}_{t}$ is a $N$-dimensional multivariate time series $X_{i,t}$ for $i\in V$ and times $t=1,\ldots,T$ for some integer $T>0$. We use the notation $i\leftrightsquigarrow j$, if $i\in V$ is directly connected to $j\in V$. Given a subset of nodes $A\subset V$, then the neighbourhood set of $A$ is defined by
@@ 1.2 GNAR models
[…]
@@
The components are the GNAR model in (3) are 1. the autoregressive parameters, $\{\alpha_{i,j}\}_{i=1,\ldots,N;j=1,\ldots p}$ explain how past values of $X_{i,\cdot}$ contribute to $X_{i,t}$. Every vertex (or node) in the graph has their own autoregressive sequence. For some data sets, it is possible to set $\alpha_{i,j}=\alpha_{j}$. In other words, a common $\{\alpha_{j}\}_{j=1,\ldots,p}$ applies to each vertex and the model is then called a global GNAR process. 2. The term $\sum_{q\in{\mathcal{N}}^{(r)}(i)}\omega_{i,q,c}^{(t)}X_{q,t-j}$ elicits a contribution from each $r$th stage neighbour, $q$, of vertex $i$, lagged by time $j$ relating to covariate type $c$. The $w^{(t)}_{i,q,c}$ are called connection weights, which are often related to the local positioning of neighbours of vertex $i$ relative to $i$ and each other. The weights could be inverse distance weights, see gnar for a definition and an example of how they are used. 3. The $s_{j}$ control the maximum number of stages of neighbours at lag $j$ for node $i$.
@@ 1.3 Features of GNAR models
Once formulated and model selection completed, GNAR models can be fit efficiently and rapidly by using existing linear modelling software. gnar and nw20 demonstrate that well-fitting GNAR-based models are highly parsimonious and also deal with missing data well.
@@ 1.4 Some univariate count time series models
We briefly review three types of popular time series models: the INAR, GAR and NAR models and some closely related ones. See cts21 for a comprehensive recent review.
The INteger-valued AutoRegressive model is based on thinning operations. The INAR(1)(1) model was introduced by inar1 and the general INAR(p)(p) model by inarp. The binomial thinning operation is defined as follows. Let $Y_{1},Y_{2},\ldots$ be a collection of independent and identically distributed Bernoulli random variables with (probability of success) parameter $q\in(0,1)$. Let $X$ be a non-negative integer random variable. The binomially-thinned random variable $Y$ is given by
@@
Let $\{\alpha_{j}\in(0,1)\}_{j=1}^{p}$, then the univariate INAR(p)(p) model for count time series, $X_{t}$ is given by
$X_{t}=\sum_{j=1}^{p}\alpha_{j}\circ X_{t-j}+\epsilon_{t}.$ (6)
Here $X_{t}\in{\mathbb{N}}_{0}$, the set of non-negative integers, and $\epsilon_{t}\in{\mathbb{N}}_{0}$ is a set of uncorrelated random variables. If the $\{\epsilon_{t}\}$ are Poisson-distributed, then $X_{t}$ is then called a Poisson-INAR(p)(p) process.
INAR(p(p) processes share many similarities with AR(p)(p) processes, e.g. the autocorrelation structure (inarp) and conditions for stationarity. However, there are differences too, such as with the conditional variance or the precise form of the autocorrelation function. Indeed, the autocorrelation of an INAR(p)(p) process has the same form as an ARMA(p,p1)(p,p-1) process, see inarp-2.
@@
The generalized autoregressive GAR(p)(p) model, without covariates, is characterised by the following ‘mean-relation’
$g(\mu_{t})=\sum_{j=1}^{p}\alpha_{j}\mathcal{A}(X_{t-j}),$ (7)
where $\mu_{t}={\mathbb{E}}(X_{t})$, $g$ is a link function, $\alpha_{j}\in{\mathbb{R}}$ and $\mathcal{A}$ is some function that modifies the autoregressive relations. GAR processes are related to generalised linear models. GAR(p)(p) models are special cases of the GARMA(p,q)(p,q) model as introduced by garma. Often, if $X_{t}\sim\text{GARMA}(p,q)$ given past history follows some exponential family, and popular choices for the conditional distribution are Poisson, binomial and gamma. Many popular nonlinear models extended from the integer-valued generalized autoregressive conditional heteroskedasticity (INGARCH) model use the same link function idea as GARMA. For example, the log-linear Poisson autoregression from loglinear or the $\operatorname{softplus}$-INGARCH model from softplusgnarch.
The nonlinear autoregressive (NAR) process nar has similarities to GAR — they both involve a link or response function as follows:
@@ 1.5 Poisson network autoregression
The first network time series model for count data was the Poisson network autoregression (PNAR) and Poisson GNAR, which are a count-valued network time series models introduced by pnar based on the network autoregression models from nvar and gnar16; gnar, respectively. The linear PNAR(p)(p) model assumes for vertex $i$ and time $t$ that $X_{i,t}\sim\operatorname{Poi}(\lambda_{i,t})$, where
@@
[…]
@@
pnar further present a count data extension of the GNAR model gnar16; gnar termed the Poisson GNAR where the conditional mean is based on GNAR components as
@@
pnar also consider a log-linear PNAR model $X_{i,t}\sim\operatorname{Poi}\{\exp(\nu_{i,t})\}$ where the linear predictor, $\nu_{i,t}$, is identical to (9) except that the $X$ terms are replaced by $\log(X+1)$ and the $\alpha$ and $\beta$ parameters can be real numbers. The +1+1 in the $\log$ term in the linear predictor is to handle zero values of $X$. Parameter estimation for both the linear and log-linear PNAR models is performed by quasi-maximised likelihood estimation (QMLE).
@@ 2.1.1 GNARI model definition
where $\circ$ denotes thinning as before, $\alpha_{i,j},\beta_{j,r}\in[0,1]$, $\epsilon_{i,t}$ is assumed to be non-negative independently distributed for each node $i$ and time $t$, and identically distributed for the same $i$, i.e. for each $i,t$ we have ${\mathbb{E}}(\epsilon_{i,t})=\lambda_{i}$ for some $\lambda_{i}\in{\mathbb{R}}^{+}$. This specification permits us to assign a different mean and variance for each node $i$, but we can choose the make the process ‘global’, i.e. $\lambda_{i}=\lambda$ as with the ‘global-$\alpha$ specification of the original GNAR processes. We assume the $\epsilon_{i,t}$ are Poisson-distributed.
GNARI models share the same limitation of not permitting negative correlations as INAR. However, they are a popular model in the regular time series case and worth study. Parameter estimation can be carried out by conditional least-squares, which we develop next.
@@ 2.1.2 GNARI conditional distribution and stationarity
We introduce the additional notation $Y^{(r)}_{i,j,t}$ to be the contribution of the $r$th-stage neighbours of node $i$ that are $j$ time steps prior to time $t$, i.e.
@@
[…]
@@ 2.1.3 GNARI process autocovariance
We now derive the autocovariance function $\Gamma(h)=\operatorname{Cov}(\mathbf{X}_{t},\mathbf{X}_{t-h})$ for (17), under stationarity. From the proof of Lemma 2.2, a GNARI($p,[s_{1},\dots,s_{p}]$) process is equivalent to the MGINAR(1) process as defined in (57). Then, by mginar Section 4, the autocovariance function for (57), $\Gamma^{\prime}(h)$, satisfies
@@
where $B$ is the variance matrix corresponding to the thinning operation $A\circ\cdot$, $\mu_{Y}=\mathbb{E}[\mathbf{Y}_{t}]$, and $\Sigma_{e}=\operatorname{Var}[\mathbf{e}_{t}]$ and $\mathbf{Y}_{t}$ is defined in (57). For GNARI, let * be entry-wise multiplication. Thus, we have
$B=\begin{bmatrix}A_{1}*(1-A_{1})&A_{2}*(1-A_{2})&\dots&A_{p-1}*(1-A_{p-1})&A_{p}*(1-A_{p})\\ 0&0&\dots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\dots&0&0\end{bmatrix}$, (21)
$\mu_{Y}=(I-A)^{-1}(\lambda_{1},\dots,\lambda_{N},0,\dots,0)^{T}$ (22)
and $\Sigma_{e}=\operatorname{diag}(\lambda_{1},\dots,\lambda_{N},0,\dots,0)$. For instance, we have
$\Gamma^{\prime}(0)=(I-A)^{-1}\{\text{diag}(B\mu_{Y})+\Sigma_{e}\}(I-A^{T})^{-1},$ (23)
which exists under stationarity.
It is easy to show that the autocovariance function for (17), $\Gamma(h+j)$, is the $[jN+1:(j+1)N,jN+1:(j+1)N]$ submatrix of $\Gamma^{\prime}(h)$. More specifically, the autocovariance function $\Gamma^{\prime}(h)$ can be written in terms of $\Gamma(h+j)$ as
@@ 2.1.4 Conditional least squares estimation
Let $\mathbf{\theta}=(\alpha_{1},\beta_{1,1},\ldots,\beta_{1,s_{1}},\ldots,\alpha_{p},\beta_{p,1},\ldots,\beta_{p,s_{p}},\lambda)^{T}$ be the parameter of interest, $\mathcal{F}_{t}$ be the $\sigma$-algebra generated by $\mathbf{X}_{t},\mathbf{X}_{t-1},\ldots$, then the conditional least squares estimator $\hat{\mathbf{\theta}}^{(n)}=\operatorname*{arg\,min}_{\theta}Q_{n}(\theta)$ minimizes
$Q_{n}(\mathbf{\theta})$ $=\sum_{t=1}^{n}||\mathbf{X}_{t}-\mathbb{E}_{\theta}(\mathbf{X}_{t}|\mathcal{F}_{t-1})||^{2}$ (25)
@@
where $S_{t,i,r,j}=\sum_{q\in\mathcal{N}^{r}_{t}(i)}w_{i,q}X_{q,t-j}$.
In practice, we use the constrained least squares algorithm described by BCL99 and implemented by the scipy.optimize.lsq_linear function from the Scipy python package, see 2020SciPy-NMeth with constraints of $[0,1]$ on the individual $\alpha$ and $\beta$ parameters. Let $\hat{\mathbf{\theta}}^{(n)}_{[0,1]}=\operatorname*{arg\,min}_{\theta}Q_{n}(\theta)$ be the constrained estimator.
@@ 2.1.5 Asymptotic properties
Let $\mathbf{X}_{t}=(X_{1,t},X_{2,t},\ldots,X_{N,t})^{T}$ is to be considered as a column vector with components that are a stationary GNARI process as defined in (12).
Definition 1.
Define $\hat{\mathbf{X}}_{t|t-1}(\mathbf{\theta})=\mathbb{E}_{\mathbf{\theta}}(\mathbf{X}_{t}|\mathcal{F}_{t-1})$ and
$f_{t|t-1}(\mathbf{\theta})=\mathbb{E}[\{\mathbf{X}_{t}-\hat{\mathbf{X}}_{t|t-1}(\mathbf{\theta})\}\{\mathbf{X}_{t}-\hat{\mathbf{X}}_{t|t-1}(\mathbf{\theta})\}^{T}|\mathcal{F}_{t-1}]$ (30)
and let $\mathbf{\theta}_{0}$ be the true value of $\mathbf{\theta}$.
We now show asymptotic consistency and that the estimator is asymptotically normal.
Proposition 1.
Assuming GNARI process stationarity, then
  1. 1.
    $\hat{\theta}^{(n)}\xrightarrow{a.s}\mathbf{\theta}_{0}.$ (31)
  2. 2.
    $n^{1/2}(\hat{\mathbf{\theta}}^{(n)}-\mathbf{\theta}_{0})\xrightarrow{}\operatorname{MVN}(0,U^{-1}RU^{-1}),$ (32)
@@
    $U=\mathbb{E}\left\{\frac{\partial\hat{\mathbf{X}}_{t|t-1}^{T}}{\partial\mathbf{\theta}}(\mathbf{\theta}_{0})\frac{\partial\hat{\mathbf{X}}_{t|t-1}}{\partial\mathbf{\theta}}(\mathbf{\theta}_{0})\right\},$ (33)
@@
    $R=\mathbb{E}\left\{\frac{\partial\hat{\mathbf{X}}_{t|t-1}^{T}}{\partial\mathbf{\theta}}(\mathbf{\theta}_{0})f_{t|t-1}(\mathbf{\theta}_{0})\frac{\partial\hat{\mathbf{X}}_{t|t-1}}{\partial\mathbf{\theta}}(\mathbf{\theta}_{0})\right\}.$ (34)
See appendix for proof.
Proposition 2.
Assuming GNARI process stationarity, then
  1. 1.
    $\hat{\theta}^{(n)}_{[0,1]}\xrightarrow{a.s}\mathbf{\theta}_{0}.$ (35)